# LRM-Zero: Training Large Reconstruction Models with Synthesized Data

### v1

@@ 6 Related works
Recent progresses advances this task with learning-based feed-forward methods [98, 88, 76, 61, 43, 48, 42]. Instead of optimization, these methods train a model from large-scale object [71, 99, 91, 23, 24] or scene [102, 69, 52] data to predict the shape directly. Besides the benefits of efficiency, these feed-forward methods can naturally support sparse-views as input (e.g., 4 to 12 input view images) because they learn data patterns from massive dataset. Some models can even go with extreme case of single-view reconstruction [98, 76, 39, 79], which needs to have data prior from realistic 3D data.
Synthetic data has been popular used in computer vision [36, 56, 31], such as in segmentation [15, 21], object detection [40], image classification [37], deblurry [73], face analysis [87], etc. In 3D vision, synthetic data is widely used because of 3D data is harder to harvest, e.g., in depth estimation [4, 67], in optical flow [26, 58, 57], in finding multi-view correspondence [85], and for improving the 3D consistency of multi-view diffusion models [90]. Specifically to reconstruction, the exploration of synthetic is mainly on specific categories, for example, for face [72], for human [32], constructions [45], or for evaluation [1, 30]. Some synthetic data are template-based [9, 33, 49, 96] and injecting human’s knowledge about the semantic. Xu et al. [94, 95] and their subsequent works [50, 7, 51, 74, 103] have leveraged procedurally synthesized data for relighting, view synthesis, and various appearance acquisition and rendering tasks. However, these methods are designed for captures under controlled lighting conditions or objects with specific materials; additionally, their data is created on a relatively small scale. We revisit their procedural data generation workflow, extending it with additional data augmentation techniques and scaling it up to train large reconstruction models.
@@ Appendix F More training stability results
We shift the data distribution of Zeroverse to include more easy data and less complex boolean-augmented shapes. In experiment 9, 10, and 11, we find that training stabilizing techniques constraints the model’s training set convergence and testing set performance. Experiment 11 further shows that by reducing the ratio of boolean-augmented complex shapes, we can stably train LRM-Zero. In our final version of Zeroverse, we aim to reduce the dataset complexity in order to achieve stable training from the data distribution side to avoid the convergence constraints that Gaussian’s scale clipping or view angle threshold entails. In our empirical observation of the training data, we notice that adding boolean augmentation to objects with many basic shapes can be over complex. In experiment 12, which has the same data distribution as our final version of Zeroverse, we adopt a more even, smooth distribution between no-aug, boolean, and wireframe augmented objects and an easier basic-shape-number distribution to favor less basic shapes per object. It performs similarly as experiment 8 and 11.

### v2

@@ 6 Related works
Recent progresses advances this task with learning-based feed-forward methods [105, 94, 80, 63, 45, 50, 44]. Instead of optimization, these methods train a model from large-scale object [73, 106, 97, 23, 24] or scene [109, 71, 54] data to predict the shape directly. Besides the benefits of efficiency, these feed-forward methods can naturally support sparse-views as input (e.g., 4 to 12 input view images) because they learn data patterns from massive dataset. Some models can even go with extreme case of single-view reconstruction [105, 80, 41, 83], which needs to have data prior from realistic 3D data. Multi-view stereo methods [78, 103, 36, 87, 26, 88] are another family of feed-forward 3D reconstruction methods, but they cannot deal with sparse-view or single-view settings since they are based on local feature matching.
Synthetic data has been popular used in computer vision [38, 58, 32, 76], such as in segmentation [15, 21], object detection [42], image classification [39], deblurry [75], face analysis [93], etc. In 3D vision, synthetic data is widely used because of 3D data is harder to harvest, e.g., in depth estimation [4, 69], in optical flow [27, 60, 59], in finding multi-view correspondence [91], and for improving the 3D consistency of multi-view diffusion models [96]. Specifically to reconstruction, the exploration of synthetic is mainly on specific categories, for example, for face [74], for human [33], constructions [47], or for evaluation [1, 31]. Some synthetic data are template-based [9, 34, 51, 102] and injecting human’s knowledge about the semantic. Xu et al. [100, 101] and their subsequent works [52, 7, 53, 77, 110] have leveraged procedurally synthesized data for relighting, view synthesis, and various appearance acquisition and rendering tasks. However, these methods are designed for captures under controlled lighting conditions or objects with specific materials; additionally, their data is created on a relatively small scale. We revisit their procedural data generation workflow, extending it with additional data augmentation techniques and scaling it up to train large reconstruction models.
@@ Appendix F More training stability results
We shift the data distribution of Zeroverse to include more easy data and less complex boolean-augmented shapes. In experiment 9, 10, and 11, we find that training stabilizing techniques constraints the model’s training set convergence and testing set performance. Experiment 11 further shows that by reducing the ratio of boolean-augmented complex shapes, we can stably train LRM-Zero. In our final version of Zeroverse, we aim to reduce the dataset complexity in order to achieve stable training from the data distribution side to avoid the convergence constraints that Gaussian’s scale clipping or view angle threshold entails. In our empirical observation of the training data, we notice that adding boolean augmentation to objects with many basic shapes can be over complex. In experiment 12, which has the same data distribution as our final version of Zeroverse, we adopt a more even, smooth distribution between no-aug, boolean, and wireframe augmented objects and an easier basic-shape-number distribution to favor less basic shapes per object. It performs similarly as experiment 8 and 11. Additionally, in experiment 13, when reducing the ratio of data with boolean difference augmentation, we do not have training instability issues with the default training hyperparameters from GS-LRM, including the 0.5 perceptual loss weight.
Appendix G Additional Experiments
Table 11: Scaling down GS-LRM’s training data size. When training on only 200K instead of 800K Objaverse data, GS-LRM’s performance drops by only 0.1 PSNR on GSO.
scaling GSO ABO
def. 1x, 80K def. 1x, 300M def. 2x, 800K
1 1x 1x 2x 29.59 0.944 0.050 28.92 0.926 0.074
2 1x 1x 0.5x 29.42 0.942 0.052 28.75 0.924 0.075
Table 12: LRM-Zero vs. LRM-Zero-Obja vs. GS-LRM at the 8-input-view, 256 resolution setting. Z means Zeroverse and O means Objaverse. The LRM-Zero (first row) and GS-LRM (second row) results are from experiment 9 in Tab. 8 and experiment 2 in Tab. 9. The LRM-Zero-Obja result (third row) is obtained by training on 800K Zeroverse data and 800K Objaverse data. While LRM-Zero-Obja outperforms LRM-Zero, it underperforms GS-LRM.
scaling GSO ABO
data Training Steps Model Size Data Size PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
def. 1x, 80K def. 1x, 300M def. 1x, 400K
Z 2x 1x 4x 31.15 0.960 0.034 29.02 0.935 0.064
O 2x 1x 2x 33.12 0.973 0.024 31.75 0.957 0.047
Z&O 2x 1x 4x 32.11 0.968 0.027 30.70 0.950 0.052
We conduct additional experiments to seek more explanations on the performance gap between LRM-Zero and GS-LRM. We perform an experiment on GS-LRM, training it on a randomly sampled subset of 200K Objaverse objects. As shown in Tab. 11, GS-LRM’s performance only drops by 0.1 PSNR on GSO. We conduct joint training on both Objaverse and Zeroverse and compare it to training only on one of the two datasets in Tab. 12. Our result shows that training on both Objaverse and Zeroverse performs better than Zeroverse only, but worse than Objaverse only. Both of these experiments likely indicate that for the single-object reconstruction task, the results start to saturate with about 200K realistic data. Since the size of Objaverse is adequate for the single object reconstruction task, and the advantages of Zeroverse are not exploited in this task. However, we believe that the advantages of Zeroverse, i.e. the data size, texture quality, controllability are more valuable when extended to other tasks, such as scene reconstruction and relighting, where data is scarse.